OpenStudy (anonymous):

if 6^y is a factor of (2^14)*(3^24) , what is the greatest possible value for y: a)7 b)8 c)14 d)18 e)36

OpenStudy (anonymous):

The way I see it.. \[ 2^{14} \cdot 3^{24} = 2_1 \cdot 2_2 \cdot 2_3 \cdot .... \cdot 2_{14} \cdot 3_1 \cdot 3_2 \cdot ... \cdot 3_{24} \\ \] While 6 is basically \(2 \cdot 3\). This means that \[ 6^y = 2^y \cdot 3^y = 2_1 \cdot 2_2 \cdot .... \cdot 2_y \cdot 3_1 \cdot 3_2 \cdot ...\cdot3_{y} \] \(2^{14} \cdot 3^{24}\) is constructed from 14 such pairs and 10 more 3's so like: \[ 2^{14} \cdot 3^{14} \cdot 3^{10} = 6^{14} \cdot 3^{10} \] Then we don't have more than 14 pairs of such components required for the '6'. Means that the largets part we can convert of this expression would go into \(6^{14}\) at max. Hopefully I'm not doing an extremely dumb mistakes, but that's how I get it. I would go with 14

OpenStudy (anonymous):

sounds good..

OpenStudy (anonymous):

The ".." tells differently, but that's good enough for me. =]

OpenStudy (anonymous):

hhhahahah. no, I believe you. Can't believe it's really that simple. you are so gooD1

OpenStudy (anonymous):

ty ty =]